OpenStudy (anonymous):

Solve for x: -2x-4>8 A. X<-6 B. x>-6 C. X<-2 D. X>-2 ⭐️help please ⭐️

OpenStudy (anonymous):

-2x-4=8 -2x=8+4 -2x=12 x=12/-2 x=-6

OpenStudy (kirbykirby):

Yes that is good, but now you have to worry about the direction of the inequality. You can keep " > " all the way through until the step x "=" 12/-2 in which case you must use x < 12/-2 You must reverse the inequality sign when you divide both sides by a negative number

OpenStudy (anonymous):

thanks :D

OpenStudy (kirbykirby):

=]